OpenStudy (anonymous):

hi i really need help with my algebra 4+42/3*2-8 order of operations i keep coming up with 20 is that right i need step by step steps

OpenStudy (pokemon23):

follow PEDMAS

OpenStudy (pokemon23):

from left to right

Directrix (directrix):

24

OpenStudy (pokemon23):

the answer is indeed 24

OpenStudy (anonymous):

thank you i know its 24 but the steps is where i jam up on

Directrix (directrix):

4 + 42/3 *2 - 8 = 4 + 14 (2) - 8 = 4 + 28 - 8 = 24

OpenStudy (pokemon23):